OpenStudy (anonymous):

What is the value of the output for the function, g(c) = 19 - c^2, if the input is 4? 11 4 3 15

OpenStudy (anonymous):

@hartnn Can you help please?

OpenStudy (whpalmer4):

\[g(c) = 19-c^2\]\[g(4) = 19-(4)^2=\]

OpenStudy (anonymous):

So the answer would be 3

OpenStudy (whpalmer4):

it would